OpenStudy (anonymous):

∫ (4x^3+6x)dx=? (By ∫, the drawing below is meant.)

OpenStudy (anonymous):



OpenStudy (rogue):

\[\int\limits_{-2}^{2} (4x^3 + 6x) dx =\left[ x^4 + 3x^2 \right]_{-2}^{2} = 28 - 28 = 0\]

OpenStudy (rogue):

The function is also odd, so you don't even have to evaluate the integral since its symmetric... http://www.wolframalpha.com/input/?i=int+4x^3%2B6x+frm+-2+to+2